# Auditing Black-Box LLM APIs with a Rank-Based Uniformity Test

### v3

@@ Abstract
As API access becomes a primary interface to large language models (LLMs), users often interact with black-box systems that offer little transparency into the deployed model. To reduce costs or maliciously alter model behaviors, API providers may discreetly serve quantized or fine-tuned variants, which can degrade performance and compromise safety. Detecting such substitutions is difficult, as users lack access to model weights and, in most cases, even output logits. To tackle this problem, we propose a rank-based uniformity test that can verify the behavioral equality of a black-box LLM to a locally deployed authentic model. Our method is accurate, query-efficient, and avoids detectable query patterns, making it robust to adversarial providers that reroute or mix responses upon the detection of testing attempts. We evaluate the approach across diverse threat scenarios, including quantization, harmful fine-tuning, jailbreak prompts, and full model substitution, showing that it consistently achieves superior statistical power over prior methods under constrained query budgets.
@@ 1 Introduction
Following Gao et al. (2025), we formulate the LLM API auditing problem as a model equality test: given query access to a target LLM API and a certified reference model of the expected configurations, the goal is to determine if the two produce statistically indistinguishable outputs on shared prompts.
@@
While several methods have been proposed for model equality testing, they each fall short in one or more of these criteria (Table 1). Existing methods include Maximum Mean Discrepancy (MMD) (Gao et al., 2025), trained text classifiers (Sun et al., 2025), identity prompting (Huang et al., 2025), and benchmark performance comparison (Chen et al., 2023a). However, Sun et al. (2025) require prohibitively many API queries; Huang et al. (2025) fail to capture model variations such as size, version, or quantization (Cai et al., 2025); and Gao et al. (2025) and Chen et al. (2023a) rely on special query distributions that can be adversarially detected and circumvented by techniques such as prompt caching (Gu et al., 2025).
@@
Benchmark (Chen et al., 2023a)
Driven by these limitations, we propose a rank-based uniformity test (RUT)—an asymmetric two-sample hypothesis test that addresses all three criteria simultaneously. In RUT, we sample one response from the target API and multiple responses from the reference model for each prompt, then compute the rank percentile of the API output within the reference distribution. If the target and reference models are identical, the percentiles should follow a uniform distribution. We detect deviations using the Cramér–von Mises test (Cramér, 1928). Our method requires only a single API call per prompt, operates effectively on real-world, user-like queries, and avoids detectable patterns that adversarial providers might exploit.
We evaluate RUT across a range of adversarial scenarios in which the API provider secretly substitutes the claimed model with an alternative. In Section 5.2, we study the case where the substitute is a quantized version of the original model. In Section 5.3, we test models augmented with a hidden jailbreaking system prompt. In Section 5.4, we examine models finetuned on instruction-following data. Finally, in Section 5.5, we consider substitution with a completely different model.
Under a fixed API query budget, we find RUT outperforms both MMD and a Kolmogorov–Smirnov test (KS) baseline across all settings. It consistently achieves higher statistical power and shows greater robustness to probabilistic substitution attacks (Figure 1). Moreover, when applied to five real-world API-deployed models (Section 5.6), our method yields detection results closely aligned with other methods and shows more robustness over string-based metrics on minor decoding mismatches.
@@
    A novel statistically-principled test for auditing LLM APIs. We propose RUT, an asymmetric two-sample-test that needs only one API call per prompt and operates effectively on natural queries, achieving query efficiency and by-deign robustness to adversarial providers.
@@
    Empirical validation across diverse threat models. We perform comprehensive experiments to validate RUT’s superior statistical power under diverse model substitution settings, including quantization, jailbreaking, supervised finetuning (SFT), full model replacement, and hardware/provider replacement.
@@ 2 Related Work
LLM fingerprinting.  Fingerprinting approaches focus on identifying LLMs by analyzing their outputs. Active fingerprinting involves injecting backdoor-like behavior (Xu et al., 2024) into an LLM via finetuning, embedding watermarks (Kirchenbauer et al., 2023; Ren et al., 2023) into a model’ s text generation process, or intentionally crafting prompts to elicit unique outputs from different LLMs (Pasquini et al., 2024). Passive fingerprinting, on the other hand, focuses on analyzing the inherent patterns in LLM-generated text (Su et al., 2023; Fu et al., 2025; Alhazbi et al., 2025). This builds on the observation that LLMs expose rich “idiosyncrasies”—distributional quirks that allow classifiers to identify a model (Sun et al., 2025). While passive fingerprinting is relevant for LLM auditing, many such methods rely on training classifiers and require substantial labeled data, making them suboptimal for auditing LLM APIs. Prior work (Cai et al., 2025) also shows they are ineffective in detecting quantized model substitution.
Auditing LLM APIs.  A growing body of work investigates whether black-box APIs faithfully serve the advertised model. The most straightforward audit is to evaluate models’ benchmark performance (Analysis, 2025; Eyuboglu et al., 2024; Chen et al., 2023b), but raw performance alone cannot expose covert substitutions or partial routing. Gao et al. (2025) formalizes the problem as Model Equality Testing and shows that a kernel-MMD test can already flag public endpoints that deviate from their open-weight checkpoints. Concurrently to our work, Cai et al. (2025) investigate the model substitution setting and show that API providers can evade detection through strategies such as model quantization, randomized substitution, and benchmark evasion. Building on these insights, we propose a method that is more robust to such attacks and extend the threat model to include a broader range of realistic scenarios, such as jailbroken or maliciously finetuned models.
@@ 3 Problem Formulation
where ${\pi}_{\text{alt}}$ is an arbitrary alternative model (e.g., quantized or adversarially finetuned version of ${\pi}_{\text{ref}}$) and $q:\mathcal{X}\!\to\![0,1]$ is an unknown routing function that may depend on the prompt and even on past interaction history (e.g., routing based on prompt caching (Gu et al., 2025)). When $q(x)$ is small but positive on a non-negligible fraction of typical prompts, the mixture in (3) significantly complicates discrimination; nevertheless, a sound detection procedure must retain statistical power under such adversarial conditions.
Detector and objectives.  A detector is a decision rule $\delta:\mathcal{D}_{N}\!\to\!\{0,1\}$ whose output $1$ means “reject $H_{0}$”. Here, $\mathcal{D}_{N}$ denotes the data collected with $N$ API queries and many local model queries. For a user-specified significance level $\alpha\in(0,1)$, the detector must satisfy the type-I error constraint $\Pr_{H_{0}}\bigl{[}\delta(\mathcal{D}_{N})=1\bigr{]}\leq\alpha$. Subject to this guarantee, the verifier seeks to maximize the statistical power
$\beta(q,{\pi}_{\text{alt}})\;=\;\Pr_{H_{1}}\!\bigl{[}\delta(\mathcal{D}_{N})=1\bigr{]},$
uniformly over a broad class of substitution policies $(q,{\pi}_{\text{alt}})$; to do so with as few costly API queries $N$ as possible; and to employ prompts that are indistinguishable from ordinary user traffic, thereby preventing the provider from detecting and rerouting “suspicious-looking” inputs.
@@ 4 Method
In this section, we introduce our rank-based uniformity test in Section 4.1 and empirically identify the most discriminative score function to maximize the detection power in Section 4.2.
@@ 4.1 Rank-Based Uniformity Test (RUT)
Thus, with an injective score function $f$, testing the uniformity of $r_{\text{tgt}}$ as defined in (4) offers a valid signal for distinguishing $\pi_{\text{tgt}}$ from $\pi_{\text{ref}}$.
@@
Discriminative score function via empirical selection.  While an ideal injective score function would guarantee sensitivity to any behavioral difference between $\pi_{\text{tgt}}$ and $\pi_{\text{ref}}$, constructing such a function for which we can calculate the CDF is generally infeasible in practice.
@@
We apply the Cramér–von Mises (CvM) test (Cramér, 1928) to assess the deviations between $\{r_{i}\}_{i=1}^{n}$ and $\text{Uniform}[0,1]$. The test evaluates the null hypothesis
@@
[…]
@@ 4.2 Score Function Selection
The RUT requires a scalar score function $f(y,x)$. To identify a function that best captures distributional differences between models, we consider five candidate functions:
  • Log-likelihood: $\log{\pi}_{\text{ref}}(y\mid x)$.
  • Token rank: the average rank of response tokens in $y$, where a token’s rank is its position in the vocabulary ordered by the ${\pi}_{\text{ref}}$’s next-token probabilities.
  • Log-rank: the average of the logarithm of the token rank.
  • Entropy: predictive entropy for $y$ under ${\pi}_{\text{ref}}(x)$.
  • Log-likelihood log-rank ratio (LRR): the ratio between log-likelihood and log-rank. (Su et al., 2023).
@@
To identify the most discriminative score function, we conduct a Monte Carlo evaluation consisting of $500$ independent trials. In each trial, we randomly select $10$ prompts from the WildChat (Zhao et al., 2024) dataset and sample $50$ completions per prompt from both ${\pi}_{\text{ref}}$ and ${\pi}_{\text{tgt}}$, using a fixed temperature of $0.5$ and a maximum length of $30$ tokens. For each candidate score function, we compute the average AUROC (Bradley, 1997) across the $10$ prompts for each trial, yielding a distribution of $500$ AUROC scores per function. The full algorithm to calculate per score function average AUROC is included in Appendix A.1. Across different model comparisons, we find that log-rank consistently yields the most separable AUROC distribution from $0.5$, indicating the strongest discriminative power. Figure 2 shows an example comparing Gemma-2-9b-it with its 4-bit quantized variant. Based on these results, we select log-rank as the scoring function for our uniformity test. Complete AUROC results are provided in Appendix A.2.
@@ 5 Experiments
In this section, we evaluate RUT across diverse model substitution scenarios, including quantization (Section 5.2), jailbreaks (Section 5.3), SFT (Section 5.4), full model replacement (Section 5.5), and real-world API providers (Section 5.6). Detection performance is compared against MMD and a KS baseline using statistical power AUC as the primary metric.
@@ 5.1 Experimental Setup
To evaluate detection performance under adversarial conditions, we simulate probabilistic substitution attacks where a fraction $q\in[0,1]$ of API queries are routed to an alternative model (e.g., quantized or fine-tuned). For each value of $q$, we estimate the statistical power, defined as the probability of correctly rejecting the null hypothesis when substitution is present. We then summarize the resulting power–substitution rate curve using the area under the curve (AUC) over $q\in[0,1]$. The AUC ranges from 0 to 1 and reflects the method’s ability to maintain high statistical power across varying levels of substitution, serving as a measure of robustness to such attacks. Higher values indicate more reliable and consistent detection performance. Figure 1 shows an example comparing Gemma-2-9b-it and its 4-bit quantized variant.
Data.  We use the WildChat dataset (Zhao et al., 2024), which contains real-world conversations between human users and ChatGPT. This dataset reflects authentic user behavior, ensuring the query distribution remains indistinguishable from typical API traffic—crucial for evasion-resistant auditing.
Baseline.  For the detection methods (Sun et al., 2025; Gao et al., 2025) that are compatible with WildChat, We primarily focus on Maximum Mean Discrepancy (MMD) (Gao et al., 2025) as the baseline, as Sun et al. (2025) is reported to fail to identify quantization (Cai et al., 2025).
We also tailor a Kolmogorov–Smirnov (KS) test baseline that uses the same information as RUT: it computes the log-rank scores from the reference model on both the target and reference model responses and applies the two-sample KS test (Darling, 1957) on these two sets of scores to estimate the $p$-value between the target and reference distributions.
@@
Across all models, we set the temperature to $0.5$ and cap generation at 30 tokens. We use vLLM (Kwon et al., 2023) on a single A6000 for all local inferences.
To estimate statistical power, we repeat each experiment over 500 Monte Carlo trials. Power is computed as the proportion of trials in which the test correctly rejects the null hypothesis at a significant level of 0.05. All statistical tests were run with Intel Xeon Gold 6230R @ 2.10GHz and 16 GB RAM per job.

### v4

@@ Abstract
As API access becomes a primary interface to large language models (LLMs), users often interact with black-box systems that offer little transparency into the deployed model. To reduce costs or maliciously alter model behaviors, API providers may discreetly serve quantized or fine-tuned variants, which can degrade performance and compromise safety. Detecting such substitutions is difficult, as users lack access to model weights and, in most cases, even output logits. To tackle this problem, we propose a Rank-based Uniformity Test (RUT) that can verify the behavioral equality of a black-box LLM to a locally deployed authentic model. Our method is accurate, query-efficient, and avoids detectable query patterns, making it robust to adversarial providers that reroute or mix responses upon the detection of testing attempts. We evaluate the approach across diverse query domains and threat scenarios, including quantization, harmful fine-tuning, jailbreak prompts, and full model substitution, showing that it consistently achieves superior detection power over prior methods under constrained query budgets. We release our code at https://github.com/xyzhu123/RUT.
@@ 1 Introduction
In this work, we focus on auditing API providers that serve open-weight models, and formulate the LLM API auditing as a model equality test following Gao et al. (2025): given query access to a target API and a certified reference model of the expected configurations, the goal is to determine if the two produce statistically indistinguishable outputs on shared prompts. The auditor is assumed to have full access to a faithful reference implementation of the claimed model, such as its decoding parameters and logit outputs.
@@
While several methods have been proposed for model equality testing, they each fall short in one or more of these criteria (Table 1). Existing methods include Maximum Mean Discrepancy (MMD) (Gao et al., 2025), trained text classifiers (Sun et al., 2025), identity prompting (Huang et al., 2025), and benchmark performance comparison (Chen et al., 2023b). However, Sun et al. (2025) require prohibitively many API queries; Huang et al. (2025) fail to capture model variations such as size, version, or quantization (Cai et al., 2025); and Gao et al. (2025) and Chen et al. (2023b) rely on special query distributions that can be adversarially detected and bypassed via prompt caching (Gu et al., 2025).
@@
Benchmark (Chen et al., 2023b)
Driven by these limitations, we propose a Rank-based Uniformity Test (RUT)—an asymmetric two-sample hypothesis test that addresses all three criteria simultaneously. In RUT, we sample one response from the target API and multiple responses from the reference model for each prompt, then compute the rank percentile of the API output within the reference distribution. If the target and reference models are identical, the percentiles should follow a uniform distribution. We detect deviations using the Cramér–von Mises test (Cramér01011928). Our method requires only a single API call per prompt, operates effectively on real-world, user-like queries, and avoids detectable patterns that adversarial providers might exploit.
We evaluate RUT across a range of adversarial scenarios in which the API provider secretly substitutes the claimed model with an alternative. In Section 5.2, we study the case where the substitute is a quantized version of the original model. In Section 5.3, we test models augmented with a hidden jailbreaking system prompt. In Section 5.4, we examine models finetuned on instruction-following data. In Section 5.5, we consider substitution with a completely different model. Finally, in Section 5.6, we test the methods on additional query distributions of math and coding.
Under a fixed API query budget, we find RUT outperforms both MMD and a Kolmogorov–Smirnov test (KS) baseline across all settings. It consistently achieves higher statistical power and shows greater robustness to probabilistic substitution attacks (Figure 1). Moreover, when applied to five real-world API-deployed models (Section 5.7), our method yields detection results closely aligned with other methods and is more robust over string-based metrics on minor decoding mismatches.
@@
    A novel test for auditing LLM APIs. We propose RUT, an asymmetric two-sample-test that needs only one API call per prompt and operates effectively on natural queries, achieving query efficiency and by-design robustness to adversarial providers.
@@
    Empirical validation across diverse threat models. We validate RUT under diverse settings, including quantization, jailbreaking, SFT, and full model replacement.
@@ 2 Related Work
LLM fingerprinting.  Fingerprinting approaches focus on identifying LLMs by analyzing their outputs. Active fingerprinting involves injecting backdoor-like behavior (Xu et al., 2024) into an LLM via finetuning, embedding watermarks (Kirchenbauer et al., 2023; Ren et al., 2023) into a model’ s text generation process, or intentionally crafting prompts to elicit unique outputs from different LLMs (Pasquini et al., 2024). Passive fingerprinting, on the other hand, focuses on analyzing the inherent patterns in LLM-generated text (Su et al., 2023; Fu et al., 2025; Alhazbi et al., 2025). This builds on the observation that LLMs expose rich “idiosyncrasies”—distributional quirks that allow classifiers to identify a model (Sun et al., 2025). While related, fingerprinting aims to authenticate the origin of the model and prevent publisher overclaim. Consequently, fingerprints are designed to remain stable under fine-tuning or deployment changes. This is opposite to our auditing objective. Prior work (Cai et al., 2025) also shows that fingerprinting methods fail to detect quantized substitutions.
Auditing LLM APIs.  A growing body of work investigates whether black-box APIs faithfully serve the advertised model. The most straightforward audit is to evaluate models’ benchmark performance (3; S. Eyuboglu, K. Goel, A. Desai, L. Chen, M. Monfort, C. Ré, and J. Zou (2024); L. Chen, M. Zaharia, and J. Y. Zou (2023a)), but raw performance alone cannot expose covert substitutions or partial routing. Gao et al. (2025) formalizes the problem as Model Equality Testing and shows that a kernel-MMD test can already flag public endpoints that deviate from their open-weight checkpoints. Concurrently to our work, Cai et al. (2025) investigate the model substitution setting and show that API providers can evade detection through strategies such as model quantization, randomized substitution, and benchmark evasion. Building on these insights, we propose a method that is more robust to such attacks and extend the threat model to include a broader range of realistic scenarios, such as jailbroken or maliciously finetuned models.
@@ 3 Problem Formulation
where ${\pi}_{\text{alt}}$ is an arbitrary alternative model (e.g., quantized or adversarially finetuned version of ${\pi}_{\text{ref}}$) and $q:\mathcal{X}\!\to\![0,1]$ is an unknown routing function that may depend on the prompt and even on past interaction history (e.g., routing based on prompt caching (Gu et al., 2025)). When $q(x)$ is small but positive on a non-negligible fraction of typical prompts, the mixture in equation 3 significantly complicates discrimination; nevertheless, a sound detection procedure must retain statistical power under such adversarial conditions.
Detector and objectives.  A detector is a decision rule $\varphi:\mathcal{D}_{N}\!\to\!\{0,1\}$ whose output $1$ means “reject $H_{0}$”. Here, $\mathcal{D}_{N}$ denotes the data collected with $N$ API queries and many local model queries. For a user-specified significance level $\alpha\in(0,1)$, the detector must satisfy the type-I error constraint $\Pr_{H_{0}}\bigl[\varphi(\mathcal{D}_{N})=1\bigr]\leq\alpha$. Subject to this guarantee, the verifier seeks to maximize the statistical power
$\beta(q,{\pi}_{\text{alt}})\;=\;\Pr_{H_{1}}\!\bigl[\varphi(\mathcal{D}_{N})=1\bigr],$
uniformly over a broad class of substitution policies $(q,{\pi}_{\text{alt}})$; to do so with as few costly API queries $N$ as possible; and to employ prompts indistinguishable from ordinary user traffic, thereby preventing the provider from detecting and rerouting “suspicious-looking” inputs.
@@ 4 Method
We introduce our rank-based uniformity test in Section 4.1 and empirically identify the most discriminative score function to maximize the detection power in Section 4.2.
@@ 4.1 Rank-Based Uniformity Test (RUT)
Thus, with an injective score function $f$, testing the uniformity of $r_{\text{tgt}}$ as defined in equation 4 offers a valid signal for distinguishing $\pi_{\text{tgt}}$ from $\pi_{\text{ref}}$.
@@
Discriminative score function via empirical selection.  While an ideal injective score function would guarantee sensitivity to any difference between $\pi_{\text{tgt}}$ and $\pi_{\text{ref}}$, constructing such a function for which we can calculate the CDF is generally infeasible in practice.
@@
We apply the Cramér–von Mises (CvM) test (Cramér01011928) to assess the deviations between $\{r_{i}\}_{i=1}^{n}$ and $\text{Uniform}[0,1]$. The test evaluates the null hypothesis
@@
$\text{$p$-value}=\mathbb{P}_{H_{0}}\left[\omega^{2}_{\text{null}}\geq\omega^{2}_{\text{obs}}\right].$
We reject $H_{0}$ and conclude that $\pi_{\text{tgt}}$ and $\pi_{\text{ref}}$ are different if $\text{$p$-value}<0.05$.
@@ 4.2 Score Function Selection
The RUT requires a scalar score function $f(y,x)$. To identify a function that best captures distributional differences between models, we consider five candidate functions. Log-likelihood: $\log{\pi}_{\text{ref}}(y\mid x)$. Token rank: the average rank of response tokens in $y$, where a token’s rank is its position in the vocabulary ordered by the ${\pi}_{\text{ref}}$’s next-token probabilities. Log-rank: the average of the logarithm of the token rank. Entropy: predictive entropy for $y$ under ${\pi}_{\text{ref}}(x)$. Log-likelihood log-rank ratio (LRR): the ratio between log-likelihood and log-rank. (Su et al., 2023).
@@
To identify the most discriminative score function, we conduct a Monte Carlo evaluation consisting of $500$ independent trials. In each trial, we randomly select $10$ prompts from the WildChat (Zhao et al., 2024) dataset and sample $50$ completions per prompt from both ${\pi}_{\text{ref}}$ and ${\pi}_{\text{tgt}}$, using a fixed temperature of $0.5$ and a maximum length of $30$ tokens. For each candidate score function, we compute the average AUROC (Bradley, 1997) across the $10$ prompts for each trial, yielding a distribution of $500$ AUROC scores per function. The full algorithm to calculate per score function average AUROC is included in Appendix A.1. Across different model comparisons, we find that log-rank consistently yields the most separable AUROC distribution from $0.5$, indicating the strongest discriminative power. Figure 2 shows an example comparing Gemma-2-9b-it with its 4-bit quantized variant. Based on the results, we select log-rank as the scoring function for RUT. Complete AUROC results are provided in Appendix A.2. We also present a formal analysis of RUT’s statistical power with the log-rank score function under adversarial perturbations in Appendix C.
@@ 5 Experiments
In this section, we evaluate RUT across diverse model substitution scenarios, including quantization (Section 5.2), jailbreaks (Section 5.3), SFT (Section 5.4), full model replacement (Section 5.5), additional query domains (Section 5.6), and real-world API providers (Section 5.7). Detection performance is compared against MMD and a KS baseline using statistical power AUC as the primary metric. We also include a case study on detecting decoding parameter mismatch in Appendix B.6 and demonstrate that RUT remains robust across models and query domains.
@@ 5.1 Experimental Setup
To evaluate detection performance under adversarial conditions, we simulate probabilistic substitution attacks where a fraction $q\in[0,1]$ of API queries are routed to an alternative model (e.g., quantized or fine-tuned). For each value of $q$, we estimate the statistical power, defined as the probability of correctly rejecting the null hypothesis when substitution is present. We then summarize the resulting power–substitution rate curve using the area under the curve (AUC) over $q\in[0,1]$. The AUC ranges from 0 to 1 and reflects the method’s ability to maintain high statistical power across varying levels of substitution, serving as a measure of robustness to such attacks. Higher values indicate more reliable and consistent detection performance. Figure 1 shows an example comparing Gemma-2-9b-it and its 4-bit quantized variant. In the following experiments, we compute 95% confidence intervals for AUCs using bootstrapping. The 95% wilson confidence intervals for statistical powers are shown as shaded area on plots in Appendix B.
Data.  We use the WildChat dataset (Zhao et al., 2024), which contains real-world conversations between human users and ChatGPT. This dataset reflects authentic user behavior, ensuring the query distribution represents typical API traffic.
Baseline.  For the detection methods (Sun et al., 2025; Gao et al., 2025) that are compatible with WildChat, We primarily focus on Maximum Mean Discrepancy (MMD) (Gao et al., 2025) as the baseline, as Sun et al. (2025) is reported to fail to identify quantization (Cai et al., 2025). We also tailor a Kolmogorov–Smirnov (KS) test baseline that uses the same information as RUT: it computes the log-rank scores from the reference model on both the target and reference model responses and applies the two-sample KS test (Darling, 1957) on these two sets of scores to estimate the $p$-value.
@@
Across all models, we set the temperature to $0.5$ and cap generation at 30 tokens. We use vLLM (Kwon et al., 2023) on a single A6000 for local inferences. Statistical power is estimated over 500 Monte Carlo trials as the proportion of trials correctly rejecting the null at $\alpha=0.05$. All tests were run with Intel Xeon Gold 6230R @ 2.10GHz and 16 GB RAM.
5.2 Detecting Quantization
We consider the setting where the API provider uses a quantized variant to substitute the claimed model. We evaluate three detection methods on quantized variants of Llama-3.2-3B-Instruct [^1] https://huggingface.co/meta-llama/Llama-3.2-3B-Instruct, Mistral-7B-Instruct-v0.3222https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.3, and Gemma-2-9B-it [^3] https://huggingface.co/google/gemma-2-9b-it, comparing each model to its 4-bit and 8-bit quantized counterparts. As shown in Table 2(a), none of the methods succeed in reliably detecting substitution for the 8-bit variants of Gemma and Mistral, where statistical power AUC remains near zero across the board. In the remaining four settings, RUT outperforms MMD and the KS baseline in three out of the four cases, demonstrating superior sensitivity to quantization-induced distributional shifts. Full statistical power curves for AUCs are provided in Appendix B.1.